OpenStudy (anonymous):

Please help! Find the moment about the x-axis of a wire of constant density δ=c that lies along the curve y=sqrt(x) from x=0 to x=1. Hint: Your answer should have a c in it!

OpenStudy (sirm3d):

 with \[\Large y=\sqrt{x}\]\[\Large M_x=\lim_{n\rightarrow \infty} \sum_{i=1}^{n}y( \underbrace{\delta\cdot\Delta s}_{\text{mass of arc}})=\int_0^1c\cdot y\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]